OpenStudy (moneybaby980):

Suzanna walked two over seven of a mile in two over five of an hour. What is her unit rate in miles per hour? two sevenths over two fifths = one and two fifths miles per hour two fifths over two sevenths = one and two fifths miles per hour two sevenths over two fifths = five over seven mile per hour two fifths over two sevenths = five over seven mile per hour

OpenStudy (anonymous):

? is the two fiths thingy the answers?

OpenStudy (anonymous):

'-'

OpenStudy (anonymous):

two fifths over two sevenths = one and two fifths miles per hour maybe

OpenStudy (anonymous):

wait lemme do this, im being lazy

OpenStudy (moneybaby980):

ok

OpenStudy (anonymous):

Yeah because I got 1 so im guessing ebcause 2/5 = 2.5 and 2/7 = 3.5 u put 2.5 over 3.5 = 2.5/3.5 which gives you one so two fifths over two sevenths = one and two fifths miles per hour

OpenStudy (anonymous):

I think the last two answers are wrong and the first 2 are somewhat right but then i saw that 2/7 cant go over 2/5 because wht wtf

OpenStudy (anonymous):

wait ye im correct it willl be 2/5 over 2/7 because 2/7 would have a larger number

OpenStudy (anonymous):

so two fifths over two sevenths = one and two fifths miles per hour

OpenStudy (anonymous):

worse case is, that maybe two fifths over two sevenths = five over seven mile per hour but I really dont see that

OpenStudy (anonymous):

GG NO RE

OpenStudy (moneybaby980):

what is GG NO RE

OpenStudy (anonymous):

GOOD GAME somthin blacks say, meaning like I did good and anyone who tried to go beyond it they cant cuz i called it out!!! HAAAAAAAAAAAAAAAAAAAAAAAA btw im black

OpenStudy (anonymous):

yea i knw, but for ex. if we were slap boxing and i got the last lick and den called out GG No Re den the fight is over and no one can get a\nother lick

OpenStudy (moneybaby980):

oh